# Sturm-Liouville problems on graphs with Robin boundary conditions

### v1

@@ 2. The Robin problems and their characteristic functions
where $q_{j}\in L_{2}(0,\ell)$ are real. Given real numbers $b_{1},\ldots,b_{p}$, we set up the Robin spectral problem on the graph $G$ by imposing the generalized (standard) Robin (-Kirchhoff) conditions at the vertices $v_{i}$, $i=1,\ldots,p$, as follows: If $v_{i}$ is a pendant vertex then for each edge $e_{j}$ incoming to (respectively, for each edge $e_{k}$ outgoing from) $v_{i}$ we set
@@ 3. Eigenvalue asymptotics
To establish the asymptotic relations (3.2)–(3.3), we will now show that for each $k$ large enough a small circle $\gamma_{k}$ centered at the $k$-th element of each of the sequences (3.5) contains exactly one zero of the dunction $\sqrt{\lambda}\mapsto f(\sqrt{\lambda})+g(\sqrt{\lambda})=\phi(\lambda,b_{1},\ldots,b_{p})$. To do that we will show that $|f(\sqrt{\lambda})|>|g(\sqrt{\lambda})|$ for all $\sqrt{\lambda}\in\gamma_{k}$ and apply Rouche’s theorem.
@@ 4. An Inverse Problem
An entire function $\omega:\sqrt{\lambda}\mapsto\omega(\sqrt{\lambda})$ of exponential type $\sigma>0$ is said to be of sine type provided the following holds:  (1) all zeros of $\omega$ lie in a horizontal strip $|\operatorname{Im}\sqrt{\lambda}|<h$, (2) for a certain fixed value $\operatorname{Im}\sqrt{\lambda}=h_{1}$ and positive constants $M_{1},M_{2}$ one has $M_{1}\leq|\omega(\operatorname{Re}\sqrt{\lambda}+ih_{1})|\leq M_{2}$ for all $\operatorname{Re}\sqrt{\lambda}\in(-\infty,\infty)$, and (3) the types of $F$ in the upper and in the lower half-planes are equal.
@@
Let $\omega_{1}$ and $\omega_{2}$ be entire functions of $\sqrt{\lambda}$ of equal exponential types $\sigma>0$. Also, we assume that $\omega_{1}$ and $\omega_{2}$ are of sine type. Furthermore, let us assume that the derivatives of $\omega_{2}$ satisfy $\omega_{2}(0)=\omega^{\prime}(0)=\ldots=\omega^{(n-1)}(0)=0$ for some $n=1,2,\ldots$ so that the function $\omega:\sqrt{\lambda}\mapsto\omega_{1}(\sqrt{\lambda})+(\sqrt{\lambda})^{-n}\omega_{2}(\sqrt{\lambda})$ is entire. Then the function $\omega$ is also of sine type.

### v2

@@ 2. The Robin problems and their characteristic functions
where $q_{j}\in L_{2}(0,\ell)$ are real. Given real numbers $b_{1},\ldots,b_{p}$, we set up the Robin spectral problem on the graph $G$ by imposing the generalized (standard) Robin (-Kirchhoff) conditions at the vertices $v_{i}$, $i=1,\ldots,p$, as follows: If $v_{i}$ is a pendant vertex then for the edge $e_{j}$ incoming to (respectively, for the edge $e_{k}$ outgoing from) $v_{i}$ we set
@@ 3. Eigenvalue asymptotics
To establish the asymptotic relations (3.2)–(3.3), we will now show that for each $k$ large enough a small circle $\gamma_{k}$ centered at the $k$-th element of each of the sequences (3.5) contains exactly one zero of the function $\sqrt{\lambda}\mapsto f(\sqrt{\lambda})+g(\sqrt{\lambda})=\phi(\lambda,b_{1},\ldots,b_{p})$. To do that we will show that $|f(\sqrt{\lambda})|>|g(\sqrt{\lambda})|$ for all $\sqrt{\lambda}\in\gamma_{k}$ and apply Rouche’s theorem.
@@ 4. An Inverse Problem
An entire function $\omega:\sqrt{\lambda}\mapsto\omega(\sqrt{\lambda})$ of exponential type $\sigma>0$ is said to be of sine type provided the following holds:  (1) all zeros of $\omega$ lie in a horizontal strip $|\operatorname{Im}\sqrt{\lambda}|<h$, (2) for a certain fixed value $\operatorname{Im}\sqrt{\lambda}=h_{1}$ and positive constants $M_{1},M_{2}$ one has $M_{1}\leq|\omega(\operatorname{Re}\sqrt{\lambda}+ih_{1})|\leq M_{2}$ for all $\operatorname{Re}\sqrt{\lambda}\in(-\infty,\infty)$, and (3) the types of $\omega$ in the upper and in the lower half-planes are equal.
@@
Let $\omega_{1}$ and $\omega_{2}$ be entire functions of $\sqrt{\lambda}$ of equal exponential types $\sigma>0$. Also, we assume that $\omega_{1}$ and $\omega_{2}$ are of sine type. Furthermore, let us assume that the derivatives of $\omega_{2}$ satisfy $\omega_{2}(0)=\omega_{2}^{\prime}(0)=\ldots=\omega_{2}^{(n-1)}(0)=0$ for some $n=1,2,\ldots$ so that the function $\omega:\sqrt{\lambda}\mapsto\omega_{1}(\sqrt{\lambda})+(\sqrt{\lambda})^{-n}\omega_{2}(\sqrt{\lambda})$ is entire. Then the function $\omega$ is also of sine type.